OpenStudy (blockcolder):

Prove that \[1+{1 \over 1!}+{1 \over 2!}+{1 \over 3!}+\cdots < 3\] WITHOUT using the fact that LHS=e.

OpenStudy (campbell_st):

its a telescoping geometric sequence and need to show that |r| <1 its \[\sum_{n=0}^{\infty} 1/n \times1/(n+1)\]

OpenStudy (blockcolder):

Doesn't that sum converge to 1?

OpenStudy (anonymous):

You can show \[\frac 1{n!} \le \frac 3{2^n}\\ \sum_{i =0}^\infty \frac 1{n!} \le \sum_{i =0}^\infty\frac 3{2^n}=3\\ \]

OpenStudy (turingtest):

@eliassaab how can you show that\[\sum_{i=0}^{\infty}{1\over n!}\le\sum_{i=0}^{\infty}{3\over 2^n}\]?

OpenStudy (turingtest):

and it should be from i=1 I think

OpenStudy (anonymous):

Prove by induction that for every n \[\frac 1{n!} \le \frac 3{2^n}\\ \]

OpenStudy (turingtest):

\[\sum_{i=1}^{\infty}{1\over n!}\le\sum_{i=1}^{\infty}{3\over 2^n}=3\]induction you say? I guess I'll try it...

OpenStudy (turingtest):

oh I got my i's and n's mixed up :S \[\sum_{n=1}^{\infty}{1\over n!}\le\sum_{n=1}^{\infty}{3\over 2^n}=3\]

OpenStudy (turingtest):

wait, LHS starts from n=0, RHS starts from n=1, eh? (if it is to =3)

OpenStudy (turingtest):

..and I am having trouble proving this by induction anyway, which is no surprise since I am pretty bad at it :(

OpenStudy (anonymous):

\[ \frac 1{1!} \le \frac 3{2^1}\\ \frac 1{2!} \le \frac 3{2^2}\\ \frac 1{3!} \le \frac 3{2^3}\\ \] Now suppose that for n> 2 \[\frac 1{n!} \le \frac 3{2^n}\\ \] then \[ \frac 1{(n+1)!} =\frac 1{n!(n+1)} \le \frac 3{2^n}\frac 1{(n+1)}\\ \le \frac 3{2^n}\frac 1{2}=\frac 3{2^{n+1}} \\ \]

OpenStudy (anonymous):

I did it for you

OpenStudy (blockcolder):

Wow. Thanks for the proof. :))

OpenStudy (turingtest):

thanks, but what am I right about the LHS starting from n=0 and the RHS starting from n=1 ? or am I losing it?

OpenStudy (turingtest):

...or does that even matter?

OpenStudy (anonymous):

@TuringTest Let me fix it for the starting n=0.

OpenStudy (turingtest):

I can see your proof is correct, but thank you for the detail :)

OpenStudy (anonymous):

You can show that \[\frac 1 {n!} \le \frac {3}{2^{n+1}} \] for \[ n \ge 4 \] So \[\sum_{i =0}^\infty \frac 1{n!} \le \sum_{n =0}^3 \frac 1{n!}+ \sum_{n =4}^\infty\frac 3{2^{n+1}}\le 3\\ \]

OpenStudy (turingtest):

and that first statement can be shown with induction the exact same way I assume, only for \(n\ge4\), right?